# Modelling vehicle and pedestrian collective dynamics: Challenges and advancements

### v1

@@ 3.3.2. Mean-field games
where the expectation value is taken on the noise of the Langevin equation. This cost includes some penalty for high velocities - see the square term $(\vec{v}_{0})^{2}$ and, with $g<0$, for high pedestrian densities $\rho$. Besides, the potential $U_{0}$, assumed large, prevents pedestrians to occupy the space under obstacles - for example here the cylinder. The last term $c_{T}$ is a terminal cost which may be useful for example if pedestrians have to reach a given location. The integral is taken over a long future ($T$ is supposed to be large), but the exponential term allows to bound the anticipation to a typical timescale $1/\gamma$.
@@
This framework allows to reproduce the anticipative behaviour of pedestrians, in particular in the case of the incoming cylinder of Fig. 5. An underlying assumption is that pedestrians can predict the future density, relying on a repeated experience. In case where pedestrians have only a limited knowledge of what is happening - for example because they cannot see some parts of the scene - it can be useful to adjust their anticipation horizon $1/\gamma$ [13].
@@ 3.4. Load balancing in pedestrian evacuation
Another experiment that reveals collective dynamics and a form of long-term anticipation behaviour is the evacuation of pedestrians under normal conditions in a geometry with multiple exits [136]. We can observe that the pedestrians spread over the different exits to balance the load and optimise the final evacuation time, whether for rectangular, curved or square geometries, as in Fig. 6. As the exits are several metres apart, load balancing requires the agent to anticipate for a few seconds to select the exit and move accordingly. This typically exceeds the short-term anticipation of the operational model and requires the inclusion of some long-term tactical decision models. This involves operational/tactical coupling in modelling the dynamics, which may operate on different time scales.
@@
Figure 6. Pedestrian evacuation experiment with four possible exits [136]. Three runs are repeated with pedestrians starting at the bottom right of the geometry (see the trajectories in the lower left panel for the run A). The pedestrian distribution over the four exits shows a load balancing with similar last exit times (see the lower right panel - the circles correspond to exits 2 and 3 while the squares correspond to exits 1 and 4; the letter identifies the run).
The tactical choice of an exit is a problem of differential games, because the problem is dynamic and the choice of an exit by a pedestrian depends on the choices of his neighbours. In addition, static distance-based shortest path algorithms prove ineffective, as the choice of exit depends on the current level of congestion on the path to the exits. Real-time pedestrian exit choices can be modelled by centralised optimisation algorithms using, for example, a Hamilton–Jacobi–Bellman equation as in a mean field game [54], non-cooperative individual behavioural rules [72, 59], or learning processes based on previous experience [29]. In general, individual estimation of pedestrian travel time can be computationally intensive. However, heuristic tactical approaches based on local density levels prove to provide realistic behaviour [59, 136]. In any case, the modelling approach requires the tactical exit choice to be dynamic and the operational and tactical modelling scales to be coupled.
@@ 4.1. Stabilisation of stop-and-go waves
Stop-and-go dynamics are currently observed in conventional road traffic flow. In addition, several recent experiments [84, 44, 22, 70, 76] show that adaptive cruise control (ACC) driver assistance systems present some stability issues, partly due to large response times of the systems [77]. In the context of driving automation, there is an important need to develop stable ACC systems that are robust to delays, latency, and stochastic perturbations in the dynamics. Indeed, many factors can disrupt the stability of the flow and lead to collective oscillations.
First experiments in the late 2010s show that stable system allows the wave to be dissipated [112]. However, this requires the system to maintain a very large time gap, which limits flow performance. Connected platooning systems, known as cooperative adaptive cruise control (CACC) systems, enable efficient automated flows, particularly for truck platooning [85, 137]. However, they rely on communication systems and are not autonomous. Various compensator modules for lag and delay in the dynamics, typically using feedback control or Smith predictor, as well as filters for noise attenuation are being developed [8, 141, 62]. However, the topic remains challenging. Recently, an experiment was conducted in the United States with up to 100 connected and autonomous vehicles to stabilise traffic flows [67], including reinforcement learning techniques to damp oscillations [56]See also https://circles-consortium.github.io/.
@@ 4.2. Developing realistic microscopic pedestrian models
The practical demand for better pedestrian dynamics models that would superside existing software suggests that one should strive for as much realism in these models as possible, hence greater sophistication. A seemingly natural conclusion would be to try and refine the description of agents, notably by better accounting for the psychological processes at play, for the motivation experienced by people and their emotions. Indeed, using the pedestrian flow at a bottleneck as an illustration, it has been shown that the crowd’s level of motivation impacts how densely people pack in front of the bottleneck [1], which in turn largely controls the flow rate at not-too-high pressure [92]. In this vein, Thalmann and Musse had thus prompted the idea that the interplay between perception and emotion determines an agent’s behaviour, and should thus be accounted for [119].
Nonetheless, Corbetta et al.’s ability to reproduce the statistics of U-turns on a staircase platform with a simple scalar potential energy description [23], irrespective of the great variety of decisions that led to such U-turns in practice, demonstrates that the dynamical effect of complex psychological factors can be captured at a high level without delving into the substance of these factors.
@@ 4.3. Including socio-psychological concepts
In many situations, a natural scientific perspective on the dynamics of pedestrians is not sufficient, but must be supplemented by sociological or psychological concepts, such as, for instance, group behaviour, social identity [107, 138] or social norms [110]. This is illustrated below with two examples.
In some situations, one can observe different behaviour of individuals which can change dynamically. In [111, 1] this can be observed in the time distance plots, Fig. 6 in [1] where the paths show that some people push their way to the front while others fall back. A more detailed analysis is provided in [129] where psychologist developed an individual annotation of the behaviour of people in a crowd. These behavioural data allow to study the space-time dynamics of pushing behaviour like the relation of distance to the entrance and the probability of pushing, see [128]. The authors showed, for example, that the probability to observe pushing behaviour increases when people approach their goal, in this experiment the entrance to a concert.
@@
A further example is the behavioural diversity and dynamic as described in [110]. The authors study a entrance situation with different motivation, where queuing but also pushing could be observed. An ethogram is crated to document the diversity of behaviour and it is shown how pedestrians change their behaviour with time, see Fig. 7. The formation of a queue and its meaning as a social rule leads to a collective phenomenon for which many people work together towards a common goal. Pairwise interaction schemes such as the superposition mechanisms in force-based models corresponding to weighted averages are too simplistic and cannot consider the collective aim to form a queue. This also applies to common velocity models where the interaction with the neighbour is helpful to model the avoidance of collision but are not usable to consider the formation of an order by queuing. Furthermore, current models are unable to take into account the triggers for behavioural changes documented in [110, Fig. 7].
To describe collective interaction concepts from psychology and sociology should be integrated in modelling of the movement of agents. These could be sociological concepts such as norms and scripts [110] but also psychological concepts such as perception [140], cognition or the influence of leadership [73].
Pairwise interaction schemes, such as the superposition mechanisms in force-based models or single interaction with the neighbour with the shortest time-to-collision in certain velocity-based models, are too simplistic to account for complex behavioural changes. In fact, such mechanisms rely on intricate collective interaction schemes based on sociological and psychological cognitive aspects, and complex threshold mechanisms [43, 148]. Other aspects rely on safety components, especially for vulnerable users such as pedestrians and cyclists in mixed urban traffic [91]. The modelling of safety behaviours is based on pedestrian risk assessment that can affect the collective dynamics of a crowd and mixed traffic flow [108, 149]. In general, considerable efforts are still needed to identify the sociological, psychological and safety aspects involved and to model mechanisms capable of taking them into account [7, 118, 1, 3].
@@ 4.4. Developing generic order parameters
Another important challenge is to quantify and characterise the collective dynamics of vehicles and pedestrians using aggregated order parameters. In general, specific order parameters are defined for each of the collective dynamics. For example, stop-and-go waves can be characterised using variability indicators of vehicle speeds or gaps, while lane formation can be determined by counting the number of pedestrians in front with the same direction [94]. However, the formulation of generic order parameters remains challenging.
Lately, a modelling paradigm for particle, vehicle and pedestrian dynamics based on port-Hamiltonian systems has been developed [63, 78, 124, 35, 104]. Port-Hamiltonian systems (PHS) are a recent but already well-established modelling approach for nonlinear physical systems [132]. They decompose the dynamics in between distance-based repulsion corresponding to the Hamiltonian part with energy conservation, and some relaxation to a tactical desired velocity corresponding to the dissipative part of the system and to the input ports. Here, the Hamiltonian can be used as a generic order parameter characterising the collective dynamics [124]. In addition, the Hamiltonian can be used as a Lyapunov function to address the system stability thanks to the dissipativity inequality without the need to linearise the system. The Hamiltonian in PHS is a promising modelling approach to quantify collective dynamics in a generic way. However, the subject remains challenging and is still under development.

### v2

@@ 3.3.2. Mean-field games
where the expectation value is taken on the noise of the Langevin equation. This cost includes some penalty for high velocities - see the square term $(\vec{v}_{0})^{2}$ - and, with $g<0$, for high pedestrian densities $\rho$. Besides, the potential $U_{0}$, assumed large, prevents pedestrians to occupy the space under obstacles - for example here the cylinder. The last term $c_{T}$ is a terminal cost which may be useful for example if pedestrians have to reach a given location. The integral is taken over a long future ($T$ is supposed to be large), but the exponential term allows to bound the anticipation to a typical timescale $1/\gamma$.
@@
This framework allows to reproduce the anticipative behaviour of pedestrians, in particular in the case of the incoming cylinder of Fig. 5. An underlying assumption is that pedestrians can predict the future density, relying on a repeated experience. In case where pedestrians have only a limited knowledge of what is happening - for example because they cannot see some parts of the scene - it can be useful to adjust their anticipation horizon $1/\gamma$ [butano_a_u2024a].
@@ 3.4. Load balancing in pedestrian evacuation
Another experiment that reveals collective dynamics and a form of long-term anticipation behaviour is the evacuation of pedestrians under normal conditions in a geometry with multiple exits [wagoum2017understanding]. We can observe that the pedestrians spread over the different exits to balance the load and optimise the final evacuation time, whether for rectangular, curved or square geometries, as in Fig. 6. As the exits are several metres apart, load balancing requires the agent to anticipate for a few seconds to select the exit and move accordingly. This typically exceeds the short-term anticipation of the operational model and requires the inclusion of some long-term tactical decision models. This involves operational/tactical coupling in modelling the dynamics, which may operate on different time scales.
@@
Figure 6. Pedestrian evacuation experiment with four possible exits [wagoum2017understanding]. Three runs are repeated with pedestrians starting at the bottom right of the geometry (see the trajectories in the lower left panel for the run A). The pedestrian distribution over the four exits shows a load balancing with similar last exit times (see the lower right panel - the circles correspond to exits 2 and 3 while the squares correspond to exits 1 and 4; the letter identifies the run).
The tactical choice of an exit is a problem of differential games, because the problem is dynamic and the choice of an exit by a pedestrian depends on the choices of his neighbours. In addition, static distance-based shortest path algorithms prove ineffective, as the choice of exit depends on the current level of congestion on the path to the exits. Real-time pedestrian exit choices can be modelled by centralised optimisation algorithms using, for example, a Hamilton–Jacobi–Bellman equation as in a mean field game [hoogendoorn2004pedestrian], non-cooperative individual behavioural rules [lo2006game, kemloh2012modeling], or learning processes based on previous experience [crociani2016multidestination]. In general, individual estimation of pedestrian travel time can be computationally intensive. However, heuristic tactical approaches based on local density levels prove to provide realistic behaviour [kemloh2012modeling, wagoum2017understanding]. In any case, the modelling approach requires the tactical exit choice to be dynamic and the operational and tactical modelling scales to be coupled.
@@ 4.1. Stabilisation of stop-and-go waves
Stop-and-go dynamics are currently observed in conventional road traffic flow. In addition, several recent experiments [milanes2014modeling, gunter2020commercially, ciuffo2021requiem, li2021car, makridis2021openacc] show that adaptive cruise control (ACC) driver assistance systems present some stability issues, partly due to large response times of the systems [makridis2019response]. In the context of driving automation, there is an important need to develop stable ACC systems that are robust to delays, latency, and stochastic perturbations in the dynamics. Indeed, many factors can disrupt the stability of the flow and lead to collective oscillations.
First experiments in the late 2010s show that stable system allows the wave to be dissipated [stern2018dissipation]. However, this requires the system to maintain a very large time gap, which limits flow performance. Connected platooning systems, known as cooperative adaptive cruise control (CACC) systems, enable efficient automated flows, particularly for truck platooning [milanes2013cooperative, wang2017developing]. However, they rely on communication systems and are not autonomous. Various compensator modules for lag and delay in the dynamics, typically using feedback control or Smith predictor, as well as filters for noise attenuation are being developed [bekiaris2020pde, xing2020compensation, khound2023unified]. However, the topic remains challenging. Recently, an experiment was conducted in the United States with up to 100 connected and autonomous vehicles to stabilise traffic flows [lee2024traffic], including reinforcement learning techniques to damp oscillations [jang2024reinforcement]See also https://circles-consortium.github.io/.
@@ 4.2. Developing realistic microscopic pedestrian models
The practical demand for better pedestrian dynamics models that would superside existing software suggests that one should strive for as much realism in these models as possible, hence greater sophistication. A seemingly natural conclusion would be to try and refine the description of agents, notably by better accounting for the psychological processes at play, for the motivation experienced by people and their emotions. Indeed, using the pedestrian flow at a bottleneck as an illustration, it has been shown that the crowd’s level of motivation impacts how densely people pack in front of the bottleneck [adrian2020crowds], which in turn largely controls the flow rate at not-too-high pressure [nicolas2017pedestrian]. In this vein, Thalmann and Musse had thus prompted the idea that the interplay between perception and emotion determines an agent’s behaviour, and should thus be accounted for [thalmann2012crowd].
Nonetheless, Corbetta et al.’s ability to reproduce the statistics of U-turns on a staircase platform with a simple scalar potential energy description [corbetta2017fluctuations], irrespective of the great variety of decisions that led to such U-turns in practice, demonstrates that the dynamical effect of complex psychological factors can be captured at a high level without delving into the substance of these factors.
@@ 4.3. Including socio-psychological concepts
In many situations, a natural scientific perspective on the dynamics of pedestrians is not sufficient, but must be supplemented by sociological or psychological concepts, such as, for instance, group behaviour, social identity [seitz2017parsimony, wijermans2022towards] or social norms [siebenbehavioral]. This is illustrated below with two examples.
In some situations, one can observe different behaviour of individuals which can change dynamically. In [sieben2017collective, adrian2020crowds] this can be observed in the time distance plots, Fig. 6 in [adrian2020crowds] where the paths show that some people push their way to the front while others fall back. A more detailed analysis is provided in [usten2022pushing] where psychologist developed an individual annotation of the behaviour of people in a crowd. These behavioural data allow to study the space-time dynamics of pushing behaviour like the relation of distance to the entrance and the probability of pushing, see [uesten2023exploring]. The authors showed, for example, that the probability to observe pushing behaviour increases when people approach their goal, in this experiment the entrance to a concert.
@@
A further example is the behavioural diversity and dynamic as described in [siebenbehavioral]. The authors study a entrance situation with different motivation, where queuing but also pushing could be observed. An ethogram is crated to document the diversity of behaviour and it is shown how pedestrians change their behaviour with time, see Fig. 7. The formation of a queue and its meaning as a social rule leads to a collective phenomenon for which many people work together towards a common goal. Pairwise interaction schemes such as the superposition mechanisms in force-based models corresponding to weighted averages are too simplistic and cannot consider the collective aim to form a queue. This also applies to common velocity models where the interaction with the neighbour is helpful to model the avoidance of collision but are not usable to consider the formation of an order by queuing. Furthermore, current models are unable to take into account the triggers for behavioural changes documented in [siebenbehavioral, Fig. 7].
To describe collective interaction concepts from psychology and sociology should be integrated in modelling of the movement of agents. These could be sociological concepts such as norms and scripts [siebenbehavioral] but also psychological concepts such as perception [wirth2023neighborhood], cognition or the influence of leadership [lombardi2020nonverbal].
Pairwise interaction schemes, such as the superposition mechanisms in force-based models or single interaction with the neighbour with the shortest time-to-collision in certain velocity-based models, are too simplistic to account for complex behavioural changes. In fact, such mechanisms rely on intricate collective interaction schemes based on sociological and psychological cognitive aspects, and complex threshold mechanisms [gregorj2023social, zhang2023review]. Other aspects rely on safety components, especially for vulnerable users such as pedestrians and cyclists in mixed urban traffic [ni2016evaluation]. The modelling of safety behaviours is based on pedestrian risk assessment that can affect the collective dynamics of a crowd and mixed traffic flow [shen2021pedestrian, zhang2021pedestrian]. In general, considerable efforts are still needed to identify the sociological, psychological and safety aspects involved and to model mechanisms capable of taking them into account [beermann2023connection, templeton2020modeling, adrian2020crowds, amini2022development].
@@ 4.4. Developing generic order parameters
Another important challenge is to quantify and characterise the collective dynamics of vehicles and pedestrians using aggregated order parameters. In general, specific order parameters are defined for each of the collective dynamics. For example, stop-and-go waves can be characterised using variability indicators of vehicle speeds or gaps, while lane formation can be determined by counting the number of pedestrians in front with the same direction [nowak2012quantitative]. However, the formulation of generic order parameters remains challenging.
Lately, a modelling paradigm for particle, vehicle and pedestrian dynamics based on port-Hamiltonian systems has been developed [knorn2014passivity, matei2019inferring, tordeux2023multi, ehrhardt2024collective, rudiger2024stability]. Port-Hamiltonian systems (PHS) are a recent but already well-established modelling approach for nonlinear physical systems [van2006port]. They decompose the dynamics in between distance-based repulsion corresponding to the Hamiltonian part with energy conservation, and some relaxation to a tactical desired velocity corresponding to the dissipative part of the system and to the input ports. Here, the Hamiltonian can be used as a generic order parameter characterising the collective dynamics [tordeux2023multi]. In addition, the Hamiltonian can be used as a Lyapunov function to address the system stability thanks to the dissipativity inequality without the need to linearise the system. The Hamiltonian in PHS is a promising modelling approach to quantify collective dynamics in a generic way. However, the subject remains challenging and is still under development.